OpenStudy (help_needed_cudi):

27. Show all work for finding the solution to this system of equations using substitution. y = 2x + 4 y = 4x – 2 25. You have $25. A box of candy costs $1.75. a. Write an inequality to express the relationship between the number of boxes of candy you could buy and the total amount of money you have. b. How many boxes of candy can you buy? i know the answer to B. its 14 but i really need help with these!!! please please!! i give medals!

OpenStudy (anonymous):

27. Set them equal to each other. 2x+4=4x-2 solve for x

OpenStudy (anonymous):

25. part a. if you have 25 dollars you can spend less than or equal to 25 dollars. So if each box of candy is $1.75 thats like saying 1.75x. Put them together and you get 1.75x is less than or equal to 25

OpenStudy (help_needed_cudi):

i still dont understand 27

OpenStudy (help_needed_cudi):

when i set them beside each other then what do i do?

OpenStudy (anonymous):

you solve for x. Move all x to one side of the equation so you'd have 2x+4=4x+2 and then moving the x you have 4=4x-2x+2. Move the +2 to the left side. You have 4-2=4x-2x. Finally, 2=2x. Divide by 2, x=1

OpenStudy (help_needed_cudi):

oh ok i was so confused but thank you i understand now

OpenStudy (anonymous):

Yep you just isolate the variable (in this case x) and then divide/multiply until you have an equation that looks like this: x=something

OpenStudy (help_needed_cudi):

i love when people dont just give answers and they explain it so ik how to do my other problems but thanks